OpenStudy (anonymous):

Refer to the illustration below and find the length of the unknown side. b = 56 and c = 65. Find a.

OpenStudy (anonymous):

sqrt(65^2-56^2)=33

OpenStudy (anonymous):

how do you solve this problem first multiply??

OpenStudy (anonymous):

no...use the formula a^2+b^2=c^2 put b and c ,then find a